# Assessing subhalo finders in cosmological hydrodynamical simulations

### v1

@@ Acknowledgements
This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure. VJFM acknowledges support by NWO through the Dark Universe Science Collaboration (OCENW.XL21.XL21.025). JXH acknowledges support from National Key R&D Program of China (2023YFA1607800, 2023YFA1607801), China Manned Space Project (No. CMS- CSST-2021-A03), and 111 project (No. B20019).
We thank Peter Behroozi, Pascal Elahi and Volker Springel for their useful comments and discussion during the preparation of this manuscript. This work has benefited from using swiftsimio (Borrow & Borrisov, 2020), both to handle particle and (sub)halo data, and to image (sub)haloes. We thank Jeger Broxteman and Will McDonald for providing a suitable bird-related acronym.
@@ A.1 Collisionless, time-persistent particle tracers
Figure 15: Schematic representation of how using gas as particle tracers of subhaloes may lead to the loss of satellites in HBT+. Each column shows the time evolution of a unique subhalo, whose bound component at a given time is represented by a circle. Different colours correspond to different evolutionary branches, whose TrackIDs are given by the legend. The shaded ellipses are FoF groups, coloured as their corresponding central subhalo. The crosses and pluses represent the centre of each subhalo, which is the position of its most bound tracer particle. The filled circle represents an individual particle. Initially, Tracks 0 and 1 are contained in the same group. The gas tracer of the central subhalo, Track 0, is blown out of the FoF group. This makes HBT+ reassign Track 1 as the new central, as Track 0 is now a hostless subgroup. This mixes the particles of the two existing subhaloes together, making HBT+ unable to identify the actual subhalo that Track 1 was meant to track. Since TrackId 0 retains its previous particles, the two subhaloes now share a large number of particles. Thus, in output $N+1$, both subhaloes are likely to have the same position in phase-space, but no merging checks are applied as HBT+ incorrectly thinks they are in separate FoF groups.
@@ A.2 Weighted host finding
[…]
Figure 16: Schematic representation (as in Fig. 15) of how noise in the particle distribution may lead to misidentification of the FoF group hosting a central subhalo in HBT+. Initially, there is a single subgroup (Track 0). Subsequently, its tracer particle is not part of any FoF group, and hence HBT+ makes Track 0 hostless. This is in contrast to the bulk of its particles, which remain within the FoF group. Since the FoF group has no subhalo associated to it, a new subhalo (Track 1) is created. Similar to the example shown in Fig. 15, the two subhaloes share a large number of particles.
@@ A.3 Phase-space calculations
Computing the offset in phase-space between two pairs of subhaloes requires assigning a location to each (equation 6). In HBT+, the higher mass subhalo of a given hierarchical pair uses the centre of mass position and velocity of its 10 most bound tracers. The lower mass subhalo only uses its most bound particle tracer to locate its centre in position and velocity space.
However, using only one particle to estimate the phase-space location of subhaloes can be prone to noise. Therefore, HBT-HERONS estimates the phase-space position of both subhaloes in a hierarchical pair using the same number of particles (10 most bound tracers by default) as the most massive of the pair. Only when the least massive subhalo of the pair is an orphan, i.e. it only has one tracer particle, do we revert to using one particle to estimate its position and velocity.
[…]
Figure 17: Schematic representation (as in Fig. 15) of how a stray particle originally associated to a satellite subhalo in a different FoF can lead to its duplication in HBT+. Initially, the particle is bound and associated to Track 1. In the next output, the particle is found in the FoF group whose central is Track 2. This means that the source subgroup of Track 2 acquires the particle, and it is found to be bound. However, the source subhalo of Track 1 still contains the particle in question, which can also be found to be bound. Hence, the particle is bound to two different subhaloes and appears duplicated in the catalogues. HBT-HERONS constrains particles to be part of the same FoF group as the Track they are associated to, if the particle are found in any FoF group. For this example, Track 1 would lose the particle since they are no longer located in the same FoF group.

### v3

@@ Acknowledgements
We thank Peter Behroozi, Pascal Elahi and Volker Springel for their useful comments and discussion during the preparation of this manuscript. We also thank the referee for a constructive report that improved the presentation of our work. JCH thanks Shaun Cole and Carlos Frenk for their discussions and suggestions. This work has benefited from using swiftsimio (Borrow & Borrisov, 2020), both to handle particle and (sub)halo data, and to image (sub)haloes. We thank Jeger Broxteman and Will McDonald for providing a suitable bird-related acronym.
This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure. VJFM acknowledges support by NWO through the Dark Universe Science Collaboration (OCENW.XL21.XL21.025). JXH acknowledges support from National Key R&D Program of China (2023YFA1607800, 2023YFA1607801), China Manned Space Project (No. CMS- CSST-2021-A03), and 111 project (No. B20019). YMB acknowledges support from UK Research and Innovation through a Future Leaders Fellowship (grant agreement MR/X035166/1) and financial support from the Swiss National Science Foundation (SNSF) under project 200021_213076
@@ A.1 Collisionless, time-persistent particle tracers
Figure 15: Schematic representation of how using gas as particle tracers of subhaloes may lead to the loss of satellites in HBT+. Each column shows the time evolution of a unique subhalo, whose bound component at a given time is represented by a circle. Different colours correspond to different evolutionary branches, whose TrackIDs are given by the legend. The shaded ellipses are FoF groups, coloured as their corresponding central subhalo. The crosses and pluses represent the centre of each subhalo, which is the position of its most bound tracer particle. The filled circle represents an individual particle. Initially, Tracks 0 and 1 are contained in the same group. The gas tracer of the central subhalo, Track 0, is blown out of the FoF group. This makes HBT+ reassign Track 1 as the new central, as Track 0 is now a hostless subhalo. This mixes the particles of the two existing subhaloes together, making HBT+ unable to identify the actual subhalo that Track 1 was meant to track. Since TrackId 0 retains its previous particles, the two subhaloes now share a large number of particles. Thus, in output $N+1$, both subhaloes are likely to have the same position in phase-space, but no merging checks are applied as HBT+ incorrectly thinks they are in separate FoF groups.
@@ A.2 Weighted host finding
$N$${\textstyle N+1}\upDeltat\displaystyle\upDelta$ tTrackID10
Figure 16: Schematic representation (as in Fig. 15) of how noise in the particle distribution may lead to misidentification of the FoF group hosting a central subhalo in HBT+. Initially, there is a single subhalo (Track 0). Subsequently, its tracer particle is not part of any FoF group, and hence HBT+ makes Track 0 hostless. This is in contrast to the bulk of its particles, which remain within the FoF group. Since the FoF group has no subhalo associated to it, a new subhalo (Track 1) is created. Similar to the example shown in Fig. 15, the two subhaloes share a large number of particles.
@@ A.3 Phase-space calculations
Computing the offset in phase-space between two pairs of subhaloes requires assigning a phase-space location to each (equation 6) and an associated dispersion. In HBT+, the higher mass subhalo of a given hierarchical pair uses the centre of mass position and velocity of its $N^{\rm min}_{\rm core}$ (20 by default) most bound tracers, which are also used to estimate the position and velocity dispersion. The lower mass subhalo only uses its most bound particle tracer to locate its centre in position and velocity space.
However, using only one particle to estimate the phase-space location of subhaloes can be prone to noise. Therefore, HBT-HERONS estimates the phase-space position of both subhaloes in a hierarchical pair using the same number of particles (10 most bound tracers by default) as the most massive of the pair. Only when the least massive subhalo of the pair is an orphan, i.e. it only has one tracer particle, do we revert to using one particle to estimate its position and velocity. Additionally, we apply equation (6) symmetrically between the two subhaloes under consideration. We compute two phase-space offsets, each using the dispersion measured from one of the two subhaloes, and we take the minimum value as the offset. This change removes the assumption that the most massive subhalo of the pair has the largest dispersion of the two subhaloes, which is not always the case for poorly resolved systems.
$N$${\textstyle N+1}\upDeltat\displaystyle\upDelta$ tTrackID102
Figure 17: Schematic representation (as in Fig. 15) of how a stray particle originally associated to a satellite subhalo in a different FoF can lead to its duplication in HBT+. Initially, the particle is bound and associated to Track 1. In the next output, the particle is found in the FoF group whose central is Track 2. This means that the source subhalo of Track 2 acquires the particle, and it is found to be bound. However, the source subhalo of Track 1 still contains the particle in question, which can also be found to be bound. Hence, the particle is bound to two different subhaloes and appears duplicated in the catalogues. HBT-HERONS constrains particles to be part of the same FoF group as the Track they are associated to, if the particle is found in any FoF group. For this example, Track 1 would lose the particle since they are no longer located in the same FoF group.
@@ A.7 Refinement of the centres of subhaloes
A.8 Re-attachment of gas particles
A fundamental assumption used by HBT+ is that mass can only be transferred from lower mass subhaloes to their more massive parent subhaloes, which are connected in a way that reflects their past accretion history. This assumption works well for collisionless particle species, but may not hold for collisional particles like gas. For example, a pair of closely interacting galaxies may experience some transfer of gas particles between them, which does not necessarily follow the flow of mass as assumed possible by HBT+. This situation can lead to the misassignment of gas mass (or even stellar mass if those particles subsequently become star forming) in the central regions of galaxies, as those particles are unable to change the subhalo they are associated to.
To prevent this from happening, HBT-HERONS performs a gas re-attachment step before subjecting subhaloes to unbinding. First, we find for each gas particle in a FoF group the nearest 10 tracer particles. We then check whether any of those neighbours share the same TrackId as the gas particle in question. If none of them do, it indicates the gas particle is well beyond the confines of the subhalo it is associated to, and hence that it should be re-attached to a more suitable subhalo. We identify the subhalo the gas particle should belong to based on the subhalo that the majority of the neighbours are associated to. This change effectively removes the assumption that gas transfer is always be hierarchical in nature.